Given the points P(-1,-6), Q(15,24), and M(7,9), prove that M is the midpoint of segment PQ.
Found 2 solutions by stanbon, solver91311:
Answer by stanbon(48535)

Given the points P(-1,-6), Q(15,24), and M(7,9), prove that M is the midpoint of segment PQ.
---------
The x-coordinate of the midpoint is the average of the endpoint
x-coordinates. Same for the y-coordinate of the midpoint.
---
Your problem:
(15 + -1)/2 = 14/2 = 7
(24 + -6)/2 = 18/2 = 9
---------------------------
Cheers,
Stan H.

Answer by solver91311(12126)



Use the mid-point formula:




So, demonstrate that:






John